# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ Abstract.
Let $N$ be a prime and $\phi$ be a Hecke-Maaß cuspidal newform for the Hecke congruence subgroup $\Gamma_{0}(N)$ in $\operatorname{SL}_{n}(\mathbb{R})$. Let $\Omega$ be an adelic compactum and let $\Omega_{N}$ be its projection to $\Gamma_{0}(N)\backslash\operatorname{SL}_{n}(\mathbb{R})/\operatorname{SO}(n)$. For any prime $n$, we prove sub-baseline bounds for the sup-norm of $\phi$ restricted to $\Omega_{N}$. Conditionally on GRH, we generalise this result to all $n\geq 2$. The methods involve a new reduction theory with level structure, based on generalisations of Atkin-Lehner operators.
@@
The author was supported by the Germany Excellence Strategy grant EXC-2047/1-390685813 through the Bonn International Graduate School of Mathematics and by the ERC Advanced Grant 101054336.
@@ 1. Introduction
We normalise the invariant measure on $X_{n}(N)$ so that it has volume asymptotically equal to $N^{(n-1)+o(1)}$. Now let $\phi$ be a Hecke-Maaß form on this space, that is, a square-integrable joint eigenfunction of the invariant differential operators and the unramified Hecke algebra. Assuming that $\left\lVert\phi\right\rVert_{2}=1$, the sup-norm problem asks for non-trivial bounds on $\left\lVert\phi\right\rVert_{\infty}$. Several parameters can be considered for this question, the most studied being the spectral parameter and the level.
@@ 1.1. Some history
This problem has a rich history and the first breakthrough in the eigenvalue aspect for $n=2$ was achieved by Iwaniec and Sarnak [IS95]. They prove that $\left\lVert\phi\right\rVert_{\infty}\ll_{N,\varepsilon}\lambda^{5/24+\varepsilon}$ for any $\varepsilon>0$. This is an improvement over the so-called local bound $\left\lVert\phi\right\rVert_{\infty}\ll_{N}\lambda^{1/4}$. Their method of using an amplified pre-trace formula remains one of the main tools for obtaining such non-trivial, sub-local bounds.
@@
In the level aspect, the baseline bound expected to hold is $\left\lVert\phi\right\rVert_{\infty}\ll_{\lambda,\varepsilon}N^{\varepsilon}$ for $\phi$ a newform. The first improvement for $n=2$ is due to Blomer and Holowinsky [BH10], with important refinements by Harcos and Templier [HT12], [HT13], and the current record bound $\left\lVert\phi\right\rVert_{\infty}\ll_{\lambda,\varepsilon}N^{1/4+\varepsilon}$ is due to Khayutin, Nelson and Steiner [KNS22]. These papers deal with the case of square-free level $N$, and bounds for general $N$ were achieved in [Sah17]. The fact that much of the work on this problem historically focused on square-free levels is in large part a consequence of using Atkin-Lehner operators. This aspect of the problem forms one of the main topics of this paper.
@@
Though many other variations of the problem exist, we consider now its development in higher rank, that is, for $n>2$. In the spectral aspect we only mention here a selection, namely the work of Blomer and Pohl [BP16] (for $\operatorname{Sp}_{4}$), Blomer and Maga [BM16] (for $\operatorname{SL}_{n}$), and Marshall [Mar14] (for more general Lie groups). They achieve power savings over the local bound for any $n\geq 2$, though they only consider the sup-norm of automorphic forms restricted to a fixed compact set. The implicit constants in their bounds thus depend on this set. An investigation of the global sup-norm is the topic of Blomer, Harcos and Maga’s paper [BHM20].
@@
The present article deals with the sup-norm problem in higher rank, in the level aspect. Despite the progress described above, there are very few results in this setting. The first result, due to Hu [Hu18], considers the case of prime-power levels $N=p^{c}$, where $c$ is large, with $\phi$ corresponding to a so-called minimal vector, thus not applying to newforms. These forms are more suitable for the $p$-adic methods employed by Hu. Similar to many results in the spectral aspect, the bounds are given for the sup-norm of the restriction to a fixed adelic compact set, which we explain below in a classical language before stating the main theorem in this paper.
The second result [Tom23] is due to the author of this paper and concerns automorphic forms on a different family of spaces $\Gamma\backslash\operatorname{SL}_{n}(\mathbb{R})/\operatorname{SO}(n)$, where $\Gamma$ is a subgroup coming from the units of an order in a division algebra of degree $n$. These spaces are compact and the bounds provided are global and in terms of their volume. The degree $n$ is restricted to prime numbers and results can only be extended partially to odd degrees.
@@
Moreover, the argument is based on the fact that proper subalgebras of division algebras of prime degree are automatically fields, and that zero is the only element of norm zero. The situation is decidedly different for the matrix algebra, whose orders give rise to the groups $\Gamma_{0}(N)$, and thus the methods of [Tom23] seem to be insufficient in this case.
@@ 1.2. Statement of results
In this paper, we consider Hecke-Maaß cuspidal newforms on $X_{n}(N)$ for $n\geq 2$ and $N$ prime. Let $\Omega\subset\operatorname{SL}_{n}(\mathbb{R})/\operatorname{SO}(n)$ be a fixed compact set and define
$\Omega_{N}\subset X_{n}(N)$
as the set of $z\in X_{n}(N)$ such that there is $\gamma\in\operatorname{SL}_{n}(\mathbb{Z})$ with $\gamma z\in\Omega$. It is easy to check that $\operatorname{vol}(\Omega_{N})\asymp_{\Omega}\operatorname{vol}(X_{n}(N))$. We investigate the sup-norm of forms restricted to $\Omega_{N}$. In adelic language, this corresponds to restricting to a fixed compact domain in $\operatorname{PGL}_{n}(\mathbb{A}_{\mathbb{Q}})$, as in [Hu18].
We prove two new results, the first of which applying to all $n\geq 2$ prime.
@@
Let $n$ and $N$ be primes. Let $\phi$ be a Hecke-Maaß cuspidal newform on $X_{n}(N)$ with spectral parameter $\mu$ and define $\Omega_{N}\subset X_{n}(N)$ with respect to a fixed compact set $\Omega\in\operatorname{SL}_{n}(\mathbb{R})/\operatorname{SO}(n)$. For large $N$, we have the bound
$\left\lVert\phi|_{\Omega_{N}}\right\rVert_{\infty}\ll_{\Omega,n,\mu,\varepsilon}N^{-\frac{1}{2n^{2}}+\varepsilon}.$
The proof involves understanding the geometric structure of the problem as well as handling rather delicate diophantine conditions. It is the latter that are not yet well enough understood in the case where $n$ is not prime. However, the geometric ideas introduced in this paper are valid in full generality and already capture a significant part of the problem. To support this claim, we present below results for all $n\geq 2$, even improving those above numerically, assuming the existence of an efficient amplifier.
@@
$\sum_{p\in\mathcal{P}}\frac{|\lambda(p)|}{p^{(n-1)/2}}\gg_{\varepsilon}L^{3/4-\varepsilon}.$ (1.1)
We prove in Lemma 3.1 that condition (1.1) is true assuming the Grand Riemann Hypothesis. It is similar to condition (1.24) in [IS95], which is checked in [Hua19] for dihedral Maaß forms and in [You18] for Eisenstein series and leads to an improved exponent in the bound of Iwaniec and Sarnak, as explained in [IS95, Remark 1.6].
@@
Let $n\geq 2$ and $N$ be a prime. Let $\phi$ be a Hecke-Maaß cuspidal newform on $X_{n}(N)$ with spectral parameter $\mu$ and define $\Omega_{N}\subset X_{n}(N)$ with respect to a fixed compact set $\Omega\in\operatorname{SL}_{n}(\mathbb{R})/\operatorname{SO}(n)$. Assuming hypothesis (1.1), we have the bound
$\left\lVert\phi|_{\Omega_{N}}\right\rVert_{\infty}\ll_{\Omega,n,\mu,\varepsilon}N^{-\frac{1}{4n}+\varepsilon}.$
@@
Considering previous work on the sup-norm problem in higher rank, the main contribution of this paper is a new counting argument, based on the reduction of the domain $\Omega_{N}$ using generalised Atkin-Lehner operators, which might be of independent interest. These arguments significantly generalise and give a new perspective on the geometric methods of Harcos and Templier [HT13], which generated many strong results for the sup-norm problem on $\operatorname{GL}(2)$ (e.g. [Blo+20], [Sah17], [Ass17]). They also seem to be fundamentally different and provide stronger results than in the spectral aspect in higher rank, where savings are inverse super-exponential in $n$ [Gil20], as opposed to our inverse polynomial savings. In any case, the methods presented here provide the first steps in tackling the level aspect in higher rank and, we believe, a useful framework for proving more general and possibly stronger results in the future.
@@ 1.3. Methods
For proving both main theorems, we employ an amplified pre-trace formula to transform the analytic issue of bounding the sup-norm into a counting problem. This is one of the most common methods of studying the sup-norm of automorphic forms and goes back to the influential paper [IS95].
@@ 1.3.1. Lattices
for each $i$, where $B(r)$ is a Euclidean ball of radius $O(r)$ around 00. Note that, since $\gamma$ is an integral matrix, $e_{i}\cdot\gamma z$ is a lattice point in $L$ determining the $i$-th row of $\gamma$. Moreover, it is important to observe that $e_{n}\cdot\gamma z$ is additionally a lattice point in the sublattice $L_{N}$. On the other side, $e_{i}\cdot z$ is simply one of the basis vectors in the basis of $L$ determined by $z$.
@@
Understanding such issues is one of the main goals of reduction theory and the geometry of numbers. However, the level structure needs to be taken into consideration and, indeed, puts serious restrictions on the prospect of success for the amplified pre-trace formula strategy. We develop a novel reduction theory with level structure in Section 5 and we describe the main ideas below.
@@ 1.3.2. Generalised Atkin-Lehner operators and reduction
In a nutshell, classical reduction theory provides a way to fit a fundamental domain for $X_{n}(1)$ inside a Siegel set (for the cusp at infinity). If $z\in\operatorname{SL}_{n}(\mathbb{R})$ lies in such a fundamental domain, its rows then provide a reduced basis for the lattice $L$, that is, a basis of vectors that are as short and as orthogonal as possible.
We also obtain in this way an interpretation of the Iwasawa coordinates of $z$ in terms of the successive minima of $L$. See Section 2.3 for more details. This is not only important for implementing the refined counting strategy described above, but also for compensating with other tools when the latter fails.
@@
For instance, solving the matrix counting problem optimally and plugging the result into the amplified pre-trace formula cannot yield sub-baseline bounds when $z$ is high enough in the cusp. One then compensates by using the Fourier expansion, which gives strong bounds in terms of the Iwasawa $y$-coordinates following from the cuspidality of our automorphic form $\phi$. This is common for many of the previous works [IS95, Lemma A.1], [HT12, Lemma 5.1], etc.
@@
In the level aspect, already $z=\operatorname{id}_{n}$ has to be treated using the Fourier bound and notice that this point certainly lies in a standard bulk $\Omega_{N}$ of $X_{n}(N)$ for $\Omega$ a compact neighbourhood of the identity. From one perspective, which we do not explicate here further, this is because of the contribution of Eisenstein series on the spectral side of the pre-trace formula. In our framework, the reason is that, even though $L$ is a perfectly balanced lattice and $z$ gives an actual orthogonal basis of shortest vectors, the sublattice $L_{N}$ is maximally imbalanced.
@@
A desirable reduction theory with level structure might thus fulfil the following. It should provide a basis for the lattice $L$ that, while perhaps not reduced, gives useful information about shortest vectors in the sublattice $L_{N}$ and about the Iwasawa coordinates, meaning the Gram-Schmidt process for the basis. It should also permit some understanding of the successive minima of both $L$ and $L_{N}$. Of course, preserving the level structure means changing bases is only allowed by matrices in $\Gamma_{0}(N)$. However, there are additional symmetries at our disposal.
@@
It was recognised early on in the treatment of the sup-norm problem in the level aspect that Atkin-Lehner operators would be useful for such reductions. It is classically not hard to see that one can fit the fundamental domain for $X_{2}(N)$, where $N$ is square-free, quotiented out by the action of these operators in a Siegel set of finite volume. This is because the Atkin-Lehner operators for $N$ square-free conjugate all cusps to the cusp at infinity. Unfortunately, for powerful levels there is a deficiency of Atkin-Lehner operators and this forms an important reason why the first and many results on the sup-norm problem are restricted to square-free levels.
@@
The first authors to connect these group theoretic facts to lattices were Harcos and Templier in [HT12, Lemma 2.2]. For example, at the level of lattices, the Fricke involution for prime levels can be understood as switching the lattices in the pair $(L,L_{N})$. Together with ideas from reduction theory, this allows us to trade imbalancedness of $L$ or $L_{N}$ for closeness of $z$ to the cusp (see loc. cit.). Effectively, when the matrix counting results are weak, the Fourier bound gets better.
@@
We then provide a different perspective on the classical Atkin-Lehner operators and show in Proposition 4.4 that the only possible generalisation in this interpretation is the Fricke involution. On the one hand, this is in contrast to the case of square-free levels in $\operatorname{PGL}(2)$, but it is also a reflection of the remarkable lack of such symmetries for powerful levels. Therefore, we first only consider the case of prime level in this paper, similar to the common restrictions in the rank-one case.
The main result of our reduction theory is given in Proposition 5.2. It satisfies the intuition from the $n=2$ case, where the bulk of the reduced fundamental domain is at $\Im(z)\asymp 1/N$. In general, there are the Iwasawa coordinates $y_{1},\ldots,y_{n-1}$ and the bulk can be found at
@@
In this region, we prove that reduced $z$ satisfy that both $L$ and $L_{N}$ are balanced in Lemma 6.1. As noted above, there is also the exceptional region $\Omega$ of the bulk, where counting results would be too weak due to imbalancedness of the lattices, but the Fourier bound suffices due to closeness to the cusp.
@@
As a historical interlude, we point out some connections of the above considerations with previous work. The Atkin-Lehner involutions were already used in the breakthrough [BH10], but balancedness of lattices was interpreted in terms of Diophantine approximation properties of the Iwasawa coordinates, using terminology from the circle method.
The language of lattices was used directly in [HT12], [HT13], and subsequent works, and lead to strong numerical improvements to the bounds. However, the counting problem is interpreted using coordinates not truly inherent to lattices. Many computations in the $\operatorname{GL}(2)$ case use, in fact, the “sporadic” symplectic nature of this group. This is not available in higher degree and the direct use of coordinates seems to be very cumbersome.
For the family of groups $\operatorname{PGL}(n)$, some ideas reminiscent of the more general strategy used here can be seen in [BHM20, Sec. 3.2]. We refer also to [Ven06], where certain aspects of the geometry of $X_{n}(N)$ are studied using lattices as well.

### v2

@@ Abstract.
[…]
@@
The author was supported by the Germany Excellence Strategy grant EXC-2047/1-390685813 through the Bonn International Graduate School of Mathematics and by the European Research Council Advanced Grant 101054336.
@@ 1. Introduction
We normalise the invariant measure on $X_{n}(N)$ so that it has volume asymptotically equivalent to $N^{(n-1)}$. Now let $\phi$ be a Hecke-Maaß form on this space, that is, a square-integrable joint eigenfunction of the invariant differential operators and the unramified Hecke algebra. Assuming that $\left\lVert\phi\right\rVert_{2}=1$, the sup-norm problem asks for non-trivial bounds on $\left\lVert\phi\right\rVert_{\infty}$. Several parameters can be considered for this question, the most studied being the spectral parameter and the level.
@@ 1.1. Some history
This problem has a rich history and the first breakthrough in the eigenvalue aspect for $n=2$ was achieved by Iwaniec and Sarnak [26]. They prove that $\left\lVert\phi\right\rVert_{\infty}\ll_{N,\varepsilon}\lambda^{5/24+\varepsilon}$ for any $\varepsilon>0$. This is an improvement over the so-called local bound $\left\lVert\phi\right\rVert_{\infty}\ll_{N}\lambda^{1/4}$. Their method of using an amplified pre-trace formula remains one of the main tools for obtaining such non-trivial, sub-local bounds.
@@
In the level aspect, the baseline bound expected to hold is $\left\lVert\phi\right\rVert_{\infty}\ll_{\lambda,\varepsilon}N^{\varepsilon}$ for $\phi$ a newform. The first improvement for $n=2$ is due to Blomer and Holowinsky [6], with important refinements by Harcos and Templier [21], [22], and the current record bound $\left\lVert\phi\right\rVert_{\infty}\ll_{\lambda,\varepsilon}N^{1/4+\varepsilon}$ is due to Khayutin, Nelson and Steiner [29]. These papers deal with the case of square-free level $N$, and bounds for general $N$ were achieved in [36]. The fact that much of the work on this problem historically focused on square-free levels is in large part a consequence of using Atkin-Lehner operators. This aspect of the problem forms one of the main topics of this paper.
@@
Though many other variations of the problem exist, we consider now its development in higher rank, that is, for $n>2$. In the spectral aspect we only mention here a selection, namely the work of Blomer and Pohl [9] (for $\operatorname{Sp}_{4}$), Blomer and Maga [8] (for $\operatorname{SL}_{n}$), and Marshall [32] (for more general Lie groups). They achieve power savings over the local bound for any $n\geq 2$, though they only consider the sup-norm of automorphic forms restricted to a fixed compact set. The implicit constants in their bounds thus depend on this set, which is necessary by a result of Brumley–Templier [12] about large values in the cusp. An investigation of the global sup-norm is the topic of Blomer, Harcos and Maga’s paper [4].
@@
The present article deals with the sup-norm problem in higher rank, in the level aspect. Despite the progress described above, there are very few results in this setting. The first result, due to Hu [24], considers the case of prime-power levels $N=p^{c}$, where $c$ is large, with $\phi$ corresponding to a so-called minimal vector, thus not applying to newforms. These forms are more suitable for the $p$-adic methods employed by Hu. Similar to many results in the spectral aspect, the bounds are given for the sup-norm of the restriction to a fixed adelic compact set.
[…]
@@
Moreover, the argument is based on the fact that proper subalgebras of division algebras of prime degree are automatically fields, and that zero is the only element of norm zero. The situation is decidedly different for the matrix algebra, whose orders give rise to the groups $\Gamma_{0}(N)$, and thus the methods of [38] seem to be insufficient in this case.
@@ 1.2. Statement of results
In this paper, we consider Hecke-Maaß cuspidal newforms on $X_{n}(N)$ for $n\geq 2$ and $N$ prime. Motivated by [12] and the necessity of doing so in the spectral aspect, we truncate cuspidal regions and focus on the “bulk” of $X_{n}(N)$, which we construct as follows.
Let $\Omega\subset X_{n}(1)$ be a fixed compact set and define the preliminary set $\Omega_{N}^{\natural}\subset X_{n}(N)$ as the preimage of $\Omega$ under the natural projection $X_{n}(N)\to X_{n}(1)$. It is easy to see that $\operatorname{vol}(\Omega_{N}^{\natural})\asymp_{\Omega}\operatorname{vol}(X_{n}(N))\sim N^{n-1}$. However, some components of the preimage spread out into the cusps of $X_{n}(N)$ and need to be truncated. We use a reduction theory with level to describe this, inspired by the case $n=2$.
For this, we study a new generalisation of the classical Fricke involution on $X_{n}(N)$, denoted by $W_{N}$. We show that $\Omega_{N}^{\natural}$ can be covered by $S\cup W_{N}S$ for a Siegel set $S$ given in terms of Iwasawa coordinates by
$\{z=n(x)a(y):|x_{ij}|\leq 1/2\text{ for }1\leq i<j\leq n\text{ and }y_{1}\gg 1/N,y_{i}\gg 1\text{ for }2\leq i\leq n-1\}.$
Most of the volume of $S$ is at $y_{1}\asymp 1/N$ and $y_{i}\asymp 1$ for all $i\geq 2$.
We now further truncate $\Omega_{N}^{\natural}$ by removing a cuspidal region $C_{N}$ with significantly smaller volume
$\operatorname{vol}(C_{N})\ll N^{(n-1)(1-1/2n^{4})}$
that is given in the domain above by the additional bounds
(1.1) $y_{1}\gg N^{-1+1/2n^{4}}\text{ or }y_{i}\gg N^{1/2n^{4}}\text{ for some }i\geq 2.$
Define now the bulk as
(1.2) $\Omega_{N}=\Omega_{N}^{\natural}\setminus C_{N},$
which, volumewise, constitutes an arbitrarily large proportion of the space $X_{n}(N)$ for large enough $N$ and $\Omega$. For more details, see Section 5.
We prove two new results, the first of which applying to all prime $n\geq 2$.
@@
Let $n$ and $N$ be primes. Let $\phi$ be a Hecke-Maaß cuspidal newform on $X_{n}(N)$ with spectral parameter $\mu$ and define $\Omega_{N}\subset X_{n}(N)$ as in (1.2) with respect to a fixed compact set $\Omega\subset X_{n}(1)$. For large $N$, we have the bound
$\left\lVert\phi|_{\Omega_{N}}\right\rVert_{\infty}\ll_{\Omega,n,\mu,\varepsilon}N^{-\frac{1}{4n^{2}}+\varepsilon}.$
The proof involves understanding the geometric structure of the problem as well as handling rather delicate diophantine conditions. It is the latter that are not yet well enough understood in the case where $n$ is not prime.
Additionally, as in the case of $n=2$, bounding $\phi$ away from the bulk, e.g. on $C_{N}$, requires strong bounds from the Whittaker expansion of $\phi$. In the level aspect, these are not available in higher rank, and further investigations present serious obstacles. These remain beyond the scope of the present paper (however, see the Appendix for some remarks).
However, the geometric ideas introduced here are valid in full generality and already capture a significant part of the problem. To support this claim, we present below results for all $n\geq 2$, even improving those above numerically, assuming the existence of an efficient amplifier.
@@
(1.3) $\sum_{p\in\mathcal{P}}\frac{|\lambda(p)|}{p^{(n-1)/2}}\gg_{\varepsilon}L^{3/4-\varepsilon}.$
We prove in Lemma 3.1 that condition (1.3) is true assuming the Grand Riemann Hypothesis. It is similar to condition (1.24) in [26], which is checked in [25] for dihedral Maaß forms and in [40] for Eisenstein series and leads to an improved exponent in the bound of Iwaniec and Sarnak, as explained in [26, Remark 1.6].
@@
Let $n\geq 2$ and $N$ be a prime. Let $\phi$ be a Hecke-Maaß cuspidal newform on $X_{n}(N)$ with spectral parameter $\mu$ and define $\Omega_{N}\subset X_{n}(N)$ as in (1.2) with respect to a fixed compact set $\Omega$. Assuming hypothesis (1.3), we have the bound
$\left\lVert\phi|_{\Omega_{N}}\right\rVert_{\infty}\ll_{\Omega,n,\mu,\varepsilon}N^{-\frac{1}{8n}+\varepsilon}.$
@@
Considering previous work on the sup-norm problem in higher rank, the main contribution of this paper is a new counting argument, based on the reduction of the domain $\Omega_{N}$ using generalised Atkin-Lehner operators, which might be of independent interest. These arguments significantly generalise and give a new perspective on the geometric methods of Harcos and Templier [22], which generated many strong results for the sup-norm problem on $\operatorname{GL}(2)$ (e.g. [3], [36], [1]). They also seem to be fundamentally different and provide stronger results than in the spectral aspect in higher rank, where savings are inverse super-exponential in $n$ [19], as opposed to our inverse polynomial savings. In any case, the methods presented here provide the first steps in tackling the level aspect in higher rank and, we believe, a useful framework for proving more general and possibly stronger results in the future.
@@ 1.3. Methods
For proving both main theorems, we employ an amplified pre-trace formula to transform the analytic issue of bounding the sup-norm into a counting problem. This is one of the most common methods of studying the sup-norm of automorphic forms and goes back to the influential paper [26].
@@ 1.3.1. Lattices
for each $i$, where $B(r)$ is a Euclidean ball of radius $O(r)$ around 0. Note that, since $\gamma$ is an integral matrix, $e_{i}\cdot\gamma z$ is a lattice point in $L$ determining the $i$-th row of $\gamma$. Moreover, it is important to observe that $e_{n}\cdot\gamma z$ is additionally a lattice point in the sublattice $L_{N}$. On the other side, $e_{i}\cdot z$ is simply one of the basis vectors in the basis of $L$ determined by $z$.
@@
Understanding such issues is one of the main goals of reduction theory and the geometry of numbers. However, the level structure needs to be taken into consideration and, indeed, puts serious restrictions on the prospect of success for the amplified pre-trace formula strategy. We develop a novel reduction theory with level structure in Section 5, and we describe the main ideas below.
@@ 1.3.2. Generalised Atkin-Lehner operators and reduction
In a nutshell, classical reduction theory provides a way to fit a fundamental domain for $X_{n}(1)$ inside a Siegel set (for the cusp at infinity). If $z\in\operatorname{SL}_{n}(\mathbb{R})$ lies in such a fundamental domain, its rows then provide a reduced basis for the lattice $L$, that is, a basis of vectors that are as short and as orthogonal as possible. We also obtain in this way an interpretation of the Iwasawa coordinates of $z$ in terms of the successive minima of $L$. This is not only important for implementing the refined counting strategy described above, but also for compensating with other tools when the latter fails.
@@
For instance, solving the matrix counting problem optimally and plugging the result into the amplified pre-trace formula cannot yield sub-baseline bounds when $z$ is high enough in the cusp. One then compensates by using the Fourier expansion, which gives strong bounds in terms of the Iwasawa $y$-coordinates following from the cuspidality of our automorphic form $\phi$. This is common for many of the previous works [26, Lemma A.1], [21, Lemma 5.1], etc.
In the level aspect, already $z=\operatorname{id}_{n}$ has to be treated using the Fourier bound. Such a point lies in $\Omega_{N}^{\natural}$, for $\Omega$ a compact neighbourhood of the identity, and can be considered as lying in a cuspidal region for this problem. It is part of the set $C_{N}$ removed from $\Omega_{N}^{\natural}$ to form the bulk $\Omega_{N}$ in (1.2).
@@
From one perspective, which we do not explicate here further, the difficulty here is caused by the contribution of Eisenstein series on the spectral side of the pre-trace formula. In our framework, the reason is that, even though $L$ is a perfectly balanced lattice and $z$ gives an actual orthogonal basis of shortest vectors, the sublattice $L_{N}$ is maximally imbalanced.
@@
A desirable reduction theory with level structure might thus fulfil the following. It should provide a basis for the lattice $L$ that, while perhaps not reduced, gives useful information about the shortest vectors in the sublattice $L_{N}$ and about the Iwasawa coordinates, meaning the Gram-Schmidt process for the basis. It should also permit some understanding of the successive minima of both $L$ and $L_{N}$. Of course, preserving the level structure means changing bases is only allowed by matrices in $\Gamma_{0}(N)$. However, there are additional symmetries at our disposal.
@@
It was recognised early on in the treatment of the sup-norm problem in the level aspect that Atkin-Lehner operators would be useful for such reductions. It is classically not hard to see that one can fit the fundamental domain for $X_{2}(N)$ with respect to the action of these operators in a single Siegel set of finite volume for $N$ square-free. This is because the Atkin-Lehner operators for $N$ square-free conjugate all cusps to the cusp at infinity. Unfortunately, for powerful levels there is a deficiency of Atkin-Lehner operators and this forms an important reason why the first and many results on the sup-norm problem are restricted to square-free levels.
@@
The first authors to connect these group theoretic facts to lattices were Harcos and Templier in [21, Lemma 2.2]. For example, at the level of lattices, the Fricke involution for prime levels can be understood as switching the lattices in the pair $(L,L_{N})$. Together with ideas from reduction theory, this allows us to trade imbalancedness of $L$ or $L_{N}$ for closeness of $z$ to the cusp (see loc. cit.). Effectively, when the matrix counting results are weak, the Fourier bound should be better.
@@
We then provide a different perspective on the classical Atkin-Lehner operators and show in Corollary 4.4 that the only possible generalisation in this interpretation is the Fricke involution. On the one hand, this is in contrast to the case of square-free levels in $\operatorname{PGL}(2)$, but it is also a reflection of the remarkable lack of such symmetries for powerful levels. Therefore, we first only consider the case of prime level in this paper, similar to the common restrictions in the rank-one case.
The main result of our reduction theory is given in Proposition 5.2. It satisfies the intuition from the $n=2$ case, where the bulk of the reduced fundamental domain is at $\Im(z)\asymp 1/N$. In general, there are the Iwasawa coordinates $y_{1},\ldots,y_{n-1}$ and the bulk (in terms of volume) can be found at
@@
More generally, one can reduce $\Omega_{N}^{\natural}$ using the Fricke involution to a subset where $y_{1}\gg 1/N$ and $y_{i}\gg 1$ for $i\geq 2$. Roughly speaking, if any of these coordinates are significantly larger than in the bulk, the counting results become too weak, as is to be expected (see the previous paragraphs and the work of Harcos–Templier). Ideally, a Fourier bound would suffice to treat these cases, yet this tool is not currently available in the necessary form. For this reason, we excise the cuspidal set $C_{N}$ in the statements of the main theorems. As noted previously, this set has significantly smaller volume than $\Omega_{N}^{\natural}$.
@@
As a historical interlude, we point out some connections of the above considerations with previous work. The Atkin-Lehner involutions were already used in the breakthrough [6], but balancedness of lattices was interpreted in terms of Diophantine approximation properties of the Iwasawa coordinates, using terminology from the circle method.
The language of lattices was used directly in [21], [22], and subsequent works, and lead to strong numerical improvements to the bounds. However, the counting problem is interpreted using coordinates not truly inherent to lattices. Many computations in the $\operatorname{GL}(2)$ case use, in fact, the “sporadic” symplectic nature of this group. This is not available in higher degree and the direct use of coordinates seems to be very cumbersome.
For the family of groups $\operatorname{PGL}(n)$, some ideas reminiscent of the more general strategy used here can be seen in [4, Sec. 3.2]. We refer also to [39], where certain aspects of the geometry of $X_{n}(N)$ are studied using lattices as well.